Solve this proportion: (4x-7)/2 = 9/4
Answer by checkley79(3341)

(4x-7)/2 = 9/4 CROSS MULTIPLY
4(4X-7)=2*9
16X-28=18
16X=18+28
16X=46
X=46/16
X=2.875 ANS.
PROOF:
(4*2.875-7)/2=9/4
(11.5-7)/2=2.25
4.5/2=2.25
2.25=2.25